Given the following variables in the open economy aggregate expenditure model, autonomous consumption (C0) = 800, autonomous investment (I0) = 600, government spending (G0) = 700, export spending (X0) = 400, autonomous import spending (M0) = 700, taxes (TP) = 200, marginal propensity to consume (c1) = 0.75, marginal propensity to invest (i1) = 0.2, and marginal propensity to import (m1) = 0.15.

a. Calculate the equilibrium level of income for the open economy aggregate expenditure model. (Show all your steps carefully.)

b. If there is a decrease in the marginal propensity to consume from 0.75 to 0.65, calculate the new equilibrium level of income and the value of the multiplier. (Again, show all your steps carefully.)
1
Expert's answer
2018-12-17T06:58:11-0500

C0 = 800, I0 = 600, G0 = 700, X0 = 400, M0 = 700, T = 200, c1 = 0.75, i1 = 0.2, m1 = 0.15.

a. The equilibrium level of income for the open economy aggregate expenditure model is:

Y = C + I + G + NX = C0 + c1(Y - T) + I0 + i1*Y + G0 + X0 - (M0 + m1*Y),

Y = 800 + 0.75(Y - 200) + 600 + 0.2Y + 700 + 400 - (700 + 0.15Y),

Y = 1650 + 0.8Y,

0.2Y = 1650,

Y = 8250.

b. If there is a decrease in the marginal propensity to consume from 0.75 to 0.65, then the new equilibrium level of income is:

Y = 800 + 0.65(Y - 200) + 600 + 0.2Y + 700 + 400 - (700 + 0.15Y),

Y = 1670 + 0.7Y,

0.3Y = 1670,

Y = 5566.67.

The value of the multiplier is:

m = 1/(1 - c1 + m1) = 1/(1 - 0.65 + 0.15) = 2.